### v7

# Sums related to Euler’s totient function

Abstract.
We obtain an upper bound for the sum $\sum_{n\leq N}(a_{n}/\varphi(a_{n}))^{s}$, where $\varphi$ is Euler’s totient function, $s\in\mathbb{N}$, and $a_{1},\ldots,a_{N}$ are positive integers (not necessarily distinct) with some restrictions. As applications, for any $t>0$, we obtain an upper bound for the number of $n\in[1,N]$ such that $a_{n}/\varphi(a_{n})>t$.
Key words and phrases:
Euler’s totient function, multiplicative functions, prime numbers.
1. Introduction
Let $\varphi$ denote Euler’s totient function. We prove
Theorem 1.1.
Let $a_{1},\ldots,a_{N}$ be positive integers (not necessarily distinct), $a_{n}\leq M$ for all $1\leq n\leq N$. For $d\in\mathbb{N}$ we set
$\omega(d)=\#\{n\leq N:a_{n}\equiv 0\text{ {(mod $d$)}}\}.$
Let $\alpha\in(0,1]$, $y=\max((\log M)^{\alpha},2)$, and $\mathcal{D}$ be the collection of square-free numbers, all of whose prime divisors lie in $(1,y]$ (we note that $1\in\mathcal{D}$). Let
$\omega(n)\leq Kg(n)$
for any $n\in\mathcal{D}$, where $K>0$ is a constant (depending on $N$) and $g(n)$ is a multiplicative function. Then
$\sum_{n=1}^{N}\bigg(\frac{a_{n}}{\varphi(a_{n})}\bigg)^{s}\leq K\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$
for any $s\in\mathbb{N}$. Here $c>0$ is an absolute constant.
Theorem 1.2.
Suppose that the assumptions of Theorem 1.1 hold, $K=\gamma N$, $0<g(p)\leq c_{0}$ for any prime $p$, and
$L=\sum_{p}\frac{g(p)}{p}<\infty.$
Then there are positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$ such that
$\sum_{n=1}^{N}\bigg(\frac{a_{n}}{\varphi(a_{n})}\bigg)^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)N$ (1.1)
for any $s\in\mathbb{N}$, and
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$ (1.2)
for any $t>0$.
From Theorems 1.1 and 1.2 we obtain the following result.
Theorem 1.3.
Let $f(n)=b_{d}n^{d}+\ldots+b_{0}$ be a polynomial with integer coefficients such that $b_{d}>0$, $(b_{d},\ldots,b_{0})=1$, and $f:\mathbb{N}\to\mathbb{N}$. Then there exist positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $f$ such that for any $x\geq 1$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{n\leq x}\Big(\frac{f(n)}{\varphi(f(n))}\Big)^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)x$ (1.3)
and
$\Big\{n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))x.$ (1.4)
Theorem 1.3 extends Corollary 1.1 in [4] which showed the inequality (1.3) but with an upper bound $\textup{exp}(s\log s+Cs)x$ and without the inequality (1.4).
Let $\mathcal{L}=\{L_{1},\ldots,L_{k}\}$ be a set of $k$ distinct linear functions with integer coefficients
$L_{i}(n)=a_{i}n+b_{i},\qquad i=1,\ldots,k.$
For $L(n)=an+b$, $a,b\in\mathbb{Z}$, we define
$\Delta_{L}=|a|\prod_{i=1}^{k}|ab_{i}-ba_{i}|.$
We note that if $a_{i}=a$ for all $1\leq i\leq k$, then
$\Delta_{L}=|a|^{k+1}\prod_{i=1}^{k}|b-b_{i}|.$
Modern application of the sieve methods involves the sums
$\sum_{(a,b)\in\Omega}\frac{\Delta_{L}}{\varphi(\Delta_{L})}$
(see, for example, [3]). Here $(a,b)$ denotes a vector and $\Omega$ is a finite set in $\mathbb{Z}^{2}$. We prove
Theorem 1.4.
Let $b_{1},\ldots,b_{k}$ be distinct integers, $a\in\mathbb{N}$, $s\in\mathbb{N}$. Let $k\geq 2,$ $x\geq 3$, $(\log x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\log x$ for all $1\leq i\leq k$. We set
$f(b)=(b-b_{1})\dots(b-b_{k}).$
Then the following statements hold.
1) If $k\geq\log\log x$ or $s\leq k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\bigg(\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\bigg)^{s}\leq\Big(c\,\frac{a}{\varphi(a)}\log k\Big)^{s}\eta\log x.$ (1.5)
2) If $2\leq k<\log\log x$ and $s>k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\log x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\bigg(\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\bigg)^{s}\leq\Big(c\,\frac{a}{\varphi(a)}\log s\Big)^{s}\eta\log x.$ (1.6)
Here $c>0$ is an absolute constant.
Theorem 1.4 extends a result of Maynard ([3, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [4] which showed the inequality (1.6) but with an upper bound
$\Big(c\,\frac{a}{\varphi(a)}\log k\Big)^{s}s!\,\eta\log x.$
From Theorem 1.2 we obtain
Theorem 1.5.
Let $f(n)=b_{d}n^{d}+\ldots+b_{0}$ be a polynomial with integer coefficients such that $b_{d}>0$, $(b_{d},\ldots,b_{0})=1$, and $f:\mathbb{P}\to\mathbb{N}$. Then there exist positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $f$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big(\frac{f(p)}{\varphi(f(p))}\Big)^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)\pi(x)$ (1.7)
and
$\Big\{p\leq x:\frac{f(p)}{\varphi(f(p))}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
In particular, we can take f(x)=x1f(x)=x-1 or $f(x)=x^{2}+1$ in Theorem 1.5.
2. Notation
We reserve the letter $p$ for primes. In particular, the sum $\sum_{p\leq K}$ should be interpreted as being over all prime numbers not exceeding $K$. We denote the number of primes not exceeding $x$ by $\pi(x)$ and the number of primes not exceeding $x$ that are congruent to $a$ modulo $k$ by $\pi(x;k,a)$. Let #A\#A denote the number of elements of a finite set $A$. We write $\mathbb{Z}$ for the set of all integers, $\mathbb{N}$ for the set of all positive integers, and $\mathbb{P}$ for the set of all primes. Let $(a_{1},\ldots,a_{n})$ be the greatest common divisor of integers $a_{1},\ldots,a_{n}$ and $a_{1},\ldots,a_{n}]$ their least common multiple. For real numbers $x,$ $y$ we also use $(x,y)$ to denote the open interval and $[x,y]$ to denote the closed interval. The usage of the notation will be clear from the context.
Let $\varphi$ denote Euler’s totient function, i. e.
$\varphi(n)=\#\{1\leq m\leq n:(m,n)=1\},\quad n\in\mathbb{N}.$
We write $\nu(n)$ for the number of distinct primes dividing $n$. By $\mu(n)$ we denote the Möbius function, which is defined to be $\mu(n)=(-1)^{\nu(n)}$ if $n$ is square-free, $\mu(n)=0$ otherwise. Let $P^{+}(n)$ denote the greatest prime factor of $n$ (by convention $P^{+}(1)=1$).
By definition, we put
$\sum_{\varnothing}=0,\qquad\prod_{\varnothing}=1.$
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. By $[x]$ we denote the largest integer not exceeding $x$ and $\{x\}=x-[x]$.
For a polynomial $f(x)$ with integer coefficients, by $\rho(f,m)$ we denote the number of solutions of the congruence $f(x)\equiv 0$ (mod $m$).
3. Proof of Theorems 1.1 and 1.2
Proof of Theorem 1.1..
We need the following result.
Lemma 3.1.
There is an absolute positive constant $c$ such that if $\alpha\in(0,1]$ and $n$ is a positive integer, then
$\frac{n}{\varphi(n)}\leq\frac{c}{\alpha}\prod_{p|n:\,\,p\leq(\log n)^{\alpha}}\biggl(1+\frac{1}{p}\biggr).$
Proof.
This is [4, Lemma 3.3]. ∎
For $1\leq n\leq N$, by Lemma 3.1 we have
$\frac{a_{n}}{\varphi(a_{n})}$ $\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq(\log a_{n})^{\alpha}}\biggl(1+\frac{1}{p}\biggr)$
$\leq\frac{c}{\alpha}\prod_{p|a_{n}:\,\,p\leq y}\biggl(1+\frac{1}{p}\biggr)=\frac{c}{\alpha}\sum_{\begin{subarray}{c}d|a_{n}:\\ P^{+}(d)\leq y\end{subarray}}\frac{\mu^{2}(d)}{d}.$
Hence
$\sum_{n=1}^{N}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}$ $\leq\Big(\frac{c}{\alpha}\Big)^{s}\sum_{n=1}^{N}\sum_{\begin{subarray}{c}d_{1},\ldots,d_{s}|a_{n}:\\ P^{+}(d_{i})\leq y\end{subarray}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}$
$=\Big(\frac{c}{\alpha}\Big)^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\sum_{\begin{subarray}{c}1\leq n\leq N:\\ d_{1}|a_{n},\ldots,\,d_{s}|a_{n}\end{subarray}}1$
$=\Big(\frac{c}{\alpha}\Big)^{s}\sum_{d_{1},\ldots,d_{s}\in\mathcal{D}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}\,\omega([d_{1},\ldots,d_{s}]).$
We obtain
$\sum_{n=1}^{N}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}\leq\Big(\frac{c}{\alpha}\Big)^{s}\sum_{n\in\mathcal{D}}\omega(n)f(n),$
where
$f(n)=\sum_{\begin{subarray}{c}d_{1},\ldots,d_{s}\in\mathcal{D}:\\ [d_{1},\ldots,d_{s}]=n\end{subarray}}\frac{\mu^{2}(d_{1})\ldots\mu^{2}(d_{s})}{d_{1}\ldots d_{s}}.$
It is easy to see that $f(n)$ is supported on $\mathcal{D}$. Suppose that $m,n\in\mathcal{D}$ and $(m,n)=1$. For any $d_{1},\ldots,d_{s}\in\mathcal{D}$ such that $d_{1},\ldots,d_{s}]=mn$ there are unique $d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}}\in\mathcal{D},$ $i=1,\ldots,s,$ such that $d_{i}^{{}^{\prime}}d_{i}^{{}^{\prime\prime}}=d_{i}$, $(d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}})=1$, $d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n$, and $d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m$. Hence
$f(nm)=\sum_{\begin{subarray}{c}d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}\in\mathcal{D}:\\ [d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n\end{subarray}}\frac{\mu^{2}(d_{1}^{{}^{\prime}})\ldots\mu^{2}(d_{s}^{{}^{\prime}})}{d_{1}^{{}^{\prime}}\ldots d_{s}^{{}^{\prime}}}\sum_{\begin{subarray}{c}d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}\in\mathcal{D}:\\ [d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m\end{subarray}}\frac{\mu^{2}(d_{1}^{{}^{\prime\prime}})\ldots\mu^{2}(d_{s}^{{}^{\prime\prime}})}{d_{1}^{{}^{\prime\prime}}\ldots d_{s}^{{}^{\prime\prime}}}=f(n)f(m).$
Since $f(n)=0$ for any $n\notin\mathcal{D}$, we obtain that $f(n)$ is a multiplicative function. It is clear that
$f(p)=\sum_{k=1}^{s}\binom{s}{k}p^{-k}=(1+p^{-1})^{s}-1$
for any prime $p\leq y$.
Since $\omega(n)\leq Kg(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}$ $\leq K\Big(\frac{c}{\alpha}\Big)^{s}\sum_{n\in\mathcal{D}}f(n)g(n)=K\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+f(p)g(p)\Big)$
$=K\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big).$
Theorem 1.1 is proved. ∎
Proof of Theorem 1.2..
Since $g(p)>0$ for any prime $p$, from Theorem 1.1 we obtain
$\sum_{n=1}^{N}\bigg(\frac{a_{n}}{\varphi(a_{n})}\bigg)^{s}$ $\leq\gamma N\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$
$\leq\gamma N\Big(\frac{c}{\alpha}\Big)^{s}\prod_{p}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big).$ (3.1)
For any prime $p>s$, by the mean value theorem there is $\xi\in(0,p^{-1})$ such that
$\Big(1+\frac{1}{p}\Big)^{s}-1=s(1+\xi)^{s-1}\frac{1}{p}<s\Big(1+\frac{1}{s}\Big)^{s-1}\frac{1}{p}<\frac{es}{p}.$ (3.2)
Since $\log(1+x)\leq x$ for any $x\geq 0$, we obtain
$\prod_{p>s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$ $\leq\prod_{p>s}\Big(1+\frac{esg(p)}{p}\Big)$
$\leq\textup{exp}\Big(es\sum_{p>s}\frac{g(p)}{p}\Big)\leq\textup{exp}(eLs).$ (3.3)
If $p\leq s$, then
$1+((1+p^{-1})^{s}-1)g(p)<1+c_{0}(1+p^{-1})^{s}\leq(1+c_{0})(1+p^{-1})^{s}\leq(1+c_{0})\textup{exp}\Big(\frac{s}{p}\Big).$
By Mertens’ theorem we have
$\prod_{p\leq s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$ $\leq(1+c_{0})^{s}\textup{exp}\Big(\sum_{p\leq s}\frac{s}{p}\Big)$
$\leq(1+c_{0})^{s}\textup{exp}(s\log\log(s+2)+c_{3}s),$ (3.4)
where $c_{3}>0$ is an absolute constant. From (3.1), (3), and (3.4) we obtain
$\sum_{n=1}^{N}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}\leq\textup{exp}(s\log\log(s+2)+Cs)N,$
where $C$ is a positive constant depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$. The inequality (1.1) is proved.
Now we prove the inequality (1.2). We have
$\textup{exp}(s\log\log(s+2)+Cs)N$ $\geq\sum_{n=1}^{N}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big(\frac{a_{n}}{\varphi(a_{n})}\Big)^{s}$
$\geq t^{s}\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}1.$
We obtain
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq\textup{exp}(s\log\log(s+2)+Cs-s\log t)N.$ (3.5)
We take
$s=[\textup{exp}(te^{-(C+1)})]+1.$
Then
$s+2=\textup{exp}(te^{-(C+1)})+3-\theta,\text{ where }\theta:=\{\textup{exp}(te^{-(C+1)})\},$
and
$\log(s+2)$ $=te^{-(C+1)}+\log\Big(1+\frac{3-\theta}{\textup{exp}(te^{-(C+1)})}\Big)$
$=te^{-(C+1)}+R_{1},\qquad 0<R_{1}\leq\frac{3}{\textup{exp}(te^{-(C+1)})}.$
We get
$\log\log(s+2)$ $=\log t-(C+1)+\log\Big(1+\frac{e^{(C+1)}R_{1}}{t}\Big)$
$=\log t-(C+1)+R_{2},\qquad 0<R_{2}\leq\frac{3e^{C+1}}{t\,\textup{exp}(te^{-(C+1)})}.$
Therefore
$s\log\log(s+2)+Cs-s\log t=-s+sR_{2},$
where
$0<sR_{2}\leq\frac{3e^{C+1}(\textup{exp}(te^{-(C+1)})+1)}{t\,\textup{exp}(te^{-(C+1)})}\leq\frac{6e^{C+1}}{t}.$
We obtain
$s\log\log(s+2)+Cs-s\log t\leq-\textup{exp}(te^{-(C+1)})+\frac{6e^{C+1}}{t}\leq-\textup{exp}(te^{-(C+1)}/2),$ (3.6)
if $t\geq t_{0}$ (here $t_{0}=t_{0}(C)=t_{0}(\gamma,L,c_{0},\alpha)$ is a positive constant depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$).
We set
$c_{2}=e^{-(C+1)}/2,\qquad c_{1}=\max(\textup{exp}(\textup{exp}(c_{2}t_{0})),1).$
Then $c_{1}$ and $c_{2}$ are positive constants depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$. From (3.5) and (3.6) we obtain
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq\textup{exp}(-\textup{exp}(c_{2}t))N,$
if $t>t_{0}$. If $0<t\leq t_{0}$, then
$\#\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}$ $\leq N\leq N\,\frac{\textup{exp}(\textup{exp}(c_{2}t_{0}))}{\textup{exp}(\textup{exp}(c_{2}t))}$
$\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N.$
We obtain
$\Big\{n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big\}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$
for any $t>0$, and the inequality (1.2) is proved. This completes the proof of Theorem 1.2. ∎
4. Proof of Theorems 1.3 and 1.4
Proof of Theorem 1.3..
We set $a_{n}=f(n)$. It is clear that $N=[x]$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. We have
$y=(\log M)^{\alpha}=(d\log x+\log\tau)^{1/2}\leq(\log x)^{3/4},$ (4.1)
if $x_{0}$ is large enough. If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\log x)^{3/4})\leq\sqrt{x}<[x]=N.$ (4.2)
We need the following result.
Lemma 4.1.
Let $d$ and $m$ be positive integers. Let
$f(x)=\sum_{i=0}^{d}b_{i}x^{i},$
where $b_{0},\ldots,b_{d}$ are integers with $(b_{0},\ldots,b_{d},m)=1$. Then
$\rho(f,m)\leq cdm^{1-1/d},$
where $c>0$ is an absolute constant.
Proof.
This is [2, Theorem 2]. ∎
By Lemma 4.1 we have
$\omega(n)=\#\{k\leq N$ $:f(k)\equiv 0\text{ {(}mod $n${)}}\}\leq\rho(f,n)\Big(\frac{N}{n}+1\Big)\leq 2\rho(f,n)\,\frac{N}{n}$
$\leq(2c)dn^{1-1/d}\,\frac{N}{n}=\frac{\gamma N}{n^{1/d}},$
where $\gamma=2cd$. The function $g(n)=n^{-1/d}$ is multiplicative. Also, $g(p)=p^{-1/d}\leq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and
$L=\sum_{p}\frac{g(p)}{p}=\sum_{p}\frac{1}{p^{1+1/d}}<\infty.$
By Theorem 1.2 there exist positive constants $C$, $c_{1}$, and $c_{2}$ (depending only on $d$) such that (1.3) and (1.4) hold.
Suppose that $1\leq x<x_{0}$. For any $n\leq x$ we have
$\frac{f(n)}{\varphi(f(n))}\leq f(n)\leq\tau n^{d}\leq\tau x_{0}^{d}=A,$
where $A=A(f)>0$ is a constant depending only on $f$. Hence,
$\sum_{n\leq x}\Big(\frac{f(n)}{\varphi(f(n))}\Big)^{s}\leq A^{s}x\leq\textup{exp}(s\log\log(s+2)+Bs)x,$
where $B=\log A$ is a positive constant depending only on $f$.
We set
$\lambda(t)=\#\Big\{n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big\}.$
If $0<t\leq A$, then
$\lambda(t)\leq[x]\leq x_{0}\leq x_{0}\frac{\textup{exp}(\textup{exp}(c_{2}A))}{\textup{exp}(\textup{exp}(c_{2}t))}=\frac{b_{1}}{\textup{exp}(\textup{exp}(c_{2}t))}\leq\frac{b_{1}}{\textup{exp}(\textup{exp}(c_{2}t))}x,$
where $b_{1}=x_{0}\,\textup{exp}(\textup{exp}(c_{2}A))$ is a positive constant depending only on $f$. If $t>A$, then
$\lambda(t)=0\leq\frac{b_{1}}{\textup{exp}(\textup{exp}(c_{2}t))}x.$
We put
$\widetilde{C}=\max(C,B),\qquad\widetilde{c}_{1}=\max(c_{1},b_{1}),\qquad\widetilde{c}_{2}=c_{2}.$
Then $\widetilde{C}$, $\widetilde{c_{1}}$ and $\widetilde{c}_{2}$ are positive constants depending only on $f$, and for any $x\geq 1$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{n\leq x}\Big(\frac{f(n)}{\varphi(f(n))}\Big)^{s}\leq\textup{exp}(s\log\log(s+2)+\widetilde{C}s)x$
and
$\Big\{n\leq x:\frac{f(n)}{\varphi(f(n))}>t\Big\}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}x.$
Theorem 1.3 is proved.
Proof of Theorem 1.4..
We assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. We set
$\Omega=\{b\in\mathbb{Z}:|b|\leq\eta\log x,b\neq b_{1},\ldots,b_{k}\}.$
Since $\varphi(mn)\geq\varphi(m)\varphi(n)$ for all positive integers $m$ and $n$, and $a^{k+1}/\varphi(a^{k+1})=a/\varphi(a)$, we obtain
$S=\sum_{b\in\Omega}\bigg(\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\bigg)^{s}\leq\Big(\frac{a}{\varphi(a)}\Big)^{s}\sum_{b\in\Omega}\bigg(\frac{|f(b)|}{\varphi(|f(b)|)}\bigg)^{s}.$ (4.3)
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\log x$ for all $1\leq i\leq k$, and therefore $|f(b)|\leq(2\log x)^{k}$. Suppose that $k\geq\log\log x$. Since $n/\varphi(n)\leq c\log\log(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c(\log k+\log\log(4\log x))\leq c_{1}\log k.$
Hence,
$S\leq\Big(\frac{a}{\varphi(a)}\,c_{1}\log k\Big)^{s}\#\Omega\leq\Big(c_{2}\,\frac{a}{\varphi(a)}\log k\Big)^{s}\eta\log x,$
where $c_{2}$ is an absolute positive constant.
Suppose that $2\leq k<\log\log x$. We take $\alpha=1/4$, $M=(2\log x)^{k}$. Then
$y=(\log M)^{\alpha}\leq 2(\log\log x)^{1/2}.$
If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(4(\log\log x)^{1/2})\leq\eta\log x.$
For $n\in\mathcal{D}$, we have
$\omega(n)=\#\{b\in\Omega:f(b)\equiv 0\text{ (mod $n$)}\}$ $\leq\Big(\frac{2\eta\log x}{n}+1\Big)\prod_{p|n}\min(p,k)$
$\leq\frac{3\eta\log x}{n}\prod_{p|n}\min(p,k).$
We set $g(1)=1,$
$g(p^{\beta})=\begin{cases}\min(p,k)/p,&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
and $g(p_{1}^{\beta_{1}}\dots p_{r}^{\beta_{r}})=g(p_{1}^{\beta_{1}})\dots g(p_{r}^{\beta_{r}})$, where $p_{1},\ldots,p_{r}$ are pairwise distinct prime numbers. Then $g(n)$ is a multiplicative function and
$\omega(n)\leq(3\eta\log x)g(n)$
for any $n\in\mathcal{D}$. Since $g(p)>0$ for any prime $p$, from Theorem 1.1 we obtain
$\sum_{b\in\Omega}\bigg(\frac{|f(b)|}{\varphi(|f(b)|)}\bigg)^{s}\leq c^{s}(\eta\log x)\prod_{p\leq y}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$
$\leq c^{s}(\eta\log x)\prod_{p}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)$ (4.4)
(here $c>0$ is an absolute constant). By (4.3) and (4.4) we have
$S\leq\Big(c\,\frac{a}{\varphi(a)}\Big)^{s}\eta\log x\prod_{p}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big).$
Suppose that $s\leq k$ (we recall that $2\leq k<\log\log x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=k/p$, we obtain
$\prod_{p>k}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\prod_{p>k}\Big(1+\frac{eks}{p^{2}}\Big)\leq c_{1}^{s},$
where $c_{1}>0$ is an absolute constant. If $p\leq k$, then $g(p)=1$ and we have
$1+((1+p^{-1})^{s}-1)g(p)=(1+p^{-1})^{s}\leq\textup{exp}\Big(\frac{s}{p}\Big).$
Hence
$\prod_{p\leq k}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\textup{exp}\Big(\sum_{p\leq k}\frac{s}{p}\Big)\leq(c_{2}\log k)^{s},$
where $c_{2}>0$ is an absolute constant. The inequality (1.5) is proved.
Suppose that $s>k$ and $2\leq k<\log\log x$. Since $g(p)\leq k/p<s/p$ for any prime $p$, we obtain
$\prod_{p>s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\prod_{p>s}\Big(1+\frac{es^{2}}{p^{2}}\Big)\leq c_{3}^{s}.$
Since $g(p)\leq 1$ for any prime $p$, we have
$\prod_{p\leq s}\Big(1+((1+p^{-1})^{s}-1)g(p)\Big)\leq\textup{exp}\Big(\sum_{p\leq s}\frac{s}{p}\Big)\leq(c_{4}\log s)^{s}.$
The inequality (1.6) is proved.
Suppose that $3\leq x<x_{0}$. For any $b\in\Omega$ we have $|f(b)|\leq(2\log x)^{k}\leq A_{0}^{k}$, where $A_{0}=2\log x_{0}$, and
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c\log\log(|f(b)|+2)\leq A_{1}\log k,$
where $A_{1}>0$ is an absolute constant. Since
$\Omega\leq 2\eta\log x+1\leq 2\log x_{0}+1$
and
$\eta\log x\geq(\log x)^{1/10}\geq(\log 3)^{1/10}>1,$
by (4.3) we have
$S\leq\Big(A_{1}\frac{a}{\varphi(a)}\log k\Big)^{s}\#\Omega\leq\Big(A_{2}\frac{a}{\varphi(a)}\log k\Big)^{s}\eta\log x,$
where $A_{2}>0$ is an absolute constant. This completes the proof of Theorem 1.4.
5. Proof of Theorem 1.5
Proof.
We set $a_{p}=f(p)$. It is clear that $N=\pi(x)$. There exists a positive constant $\tau$ depending only on $f$ such that $f(p)\leq\tau p^{d}$ for any prime $p$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. By (4.1) and (4.2) we have $y=(\log M)^{\alpha}\leq(\log x)^{3/4}$ and $n\leq\sqrt{x}$ for any $n\in\mathcal{D}$.
We need the following result.
Lemma 5.1 (The Brun - Titchmarsh inequality).
If $1\leq k<x$ and $(a,k)=1$, then
$\pi(x;k,a)<\frac{3x}{\varphi(k)\log(x/k)}.$
Proof.
This is [1, Theorem 3.8]. ∎
Fix $n\in\mathcal{D}$, and let $\alpha_{1},\ldots,\alpha_{l}$ be all solutions of the congruence $f(x)\equiv 0$ (mod $n$). We have
$l=\rho(f,n)=\prod_{p|n}\rho(f,p)=\prod_{\begin{subarray}{c}p|n\\ p\nmid b_{d}\end{subarray}}\rho(f,p)\cdot\prod_{\begin{subarray}{c}p|n\\ p|b_{d}\end{subarray}}\rho(f,p)=A\cdot B.$
For any prime $p\nmid b_{d}$, by Lagrange’s theorem we have $\rho(f,p)\leq\min(p,d)$. Trivially $\rho(f,p)\leq p$ for any prime $p$. Hence,
$A\leq\prod_{\begin{subarray}{c}p|n\\ p\nmid b_{d}\end{subarray}}\min(p,d)\leq\prod_{p|n}\min(p,d),\qquad B\leq\prod_{\begin{subarray}{c}p|n\\ p|b_{d}\end{subarray}}p\leq\prod_{p|b_{d}}p=c(f).$
We obtain
$l\leq c(f)\prod_{p|n}\min(p,d),$ (5.1)
where $c(f)>0$ is a constant depending only on $f$.
We have
$\omega(n)=\#\{p\leq x:f(p)\equiv 0\text{ (mod $n$)}\}=\sum_{i=1}^{l}\pi(x;n,\alpha_{i}).$ (5.2)
Fix $i\in\{1,\ldots,l\}$ and suppose that $(\alpha_{i},n)=1$. We have
$\log\Big(\frac{x}{n}\Big)\geq\log(\sqrt{x})=\frac{1}{2}\log x.$
Applying Lemma 5.1, we obtain
$\pi(x;n,\alpha_{i})\leq\frac{6x}{\varphi(n)\log x}.$ (5.3)
Suppose that $(\alpha_{i},n)>1$. Hence,
$\pi(x;n,\alpha_{i})\leq 1\leq\frac{6x}{\varphi(n)\log x},$
since
$\frac{6x}{\varphi(n)\log x}\geq\frac{6x}{n\log x}\geq\frac{6\sqrt{x}}{\log x}>1,$
if $x_{0}$ is large enough. We see that the inequality (5.3) holds in both cases $(\alpha_{i},n)=1$ and $(\alpha_{i},n)>1$. From (5.1) – (5.3) we get
$\omega(n)\leq\frac{6xl}{\varphi(n)\log x}\leq\frac{\gamma\pi(x)}{\varphi(n)}\prod_{p|n}\min(p,d),$
where $\gamma=\gamma(f)>0$ is a constant depending only on $f$.
We set $g(1)=1,$
$g(p^{\beta})=\begin{cases}\min(p,d)/(p-1),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
and $g(p_{1}^{\beta_{1}}\dots p_{r}^{\beta_{r}})=g(p_{1}^{\beta_{1}})\dots g(p_{r}^{\beta_{r}})$, where $p_{1},\ldots,p_{r}$ are pairwise distinct prime numbers. Then $g(n)$ is a multiplicative function and
$\omega(n)\leq\gamma\pi(x)g(n)$
for any $n\in\mathcal{D}$.
Since
$g(p)=\frac{\min(p,d)}{p-1}\leq\frac{p}{p-1}\leq 2$
for any prime $p$ and
$L=\sum_{p}\frac{g(p)}{p}=\sum_{p\leq d}\frac{1}{p-1}+\sum_{p>d}\frac{d}{p(p-1)}<\infty,$
by Theorem 1.2 there exist positive constants $C$, $c_{1}$ and $c_{2}$ depending only on $f$ such that the inequalities (1.7) and (1.8) hold.
By arguing as in the end of the proof of Theorem 1.3, we find that there exist positive constants $\widetilde{C}$, $\widetilde{c}_{1}$, and $\widetilde{c}_{2}$ depending only on $f$ such that for any $x\geq 2$, $s\in\mathbb{N}$, and $t>0$ we have
$\sum_{p\leq x}\Big(\frac{f(p)}{\varphi(f(p))}\Big)^{s}\leq\textup{exp}(s\log\log(s+2)+\widetilde{C}s)\pi(x)$
and
$\Big\{p\leq x:\frac{f(p)}{\varphi(f(p))}>t\Big\}\leq\frac{\widetilde{c}_{1}}{\textup{exp}(\textup{exp}(\widetilde{c}_{2}t))}\pi(x).$
Theorem 1.5 is proved.
6. Acknowledgements
The author is grateful to Mikhail R. Gabdullin for useful discussions and to the anonymous referee for useful comments.
This research was supported by Russian Science Foundation, grant 20-11-20203, https://rscf.ru/en/project/20-11-20203/.